\begin{thebibliography}{29}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Arcee(2024)}]{spark}
Arcee. 2024.
\newblock \href {https://huggingface.co/arcee-ai/Llama-Spark} {Llama spark model card}.

\bibitem[{Craswell et~al.(2020)Craswell, Mitra, Yilmaz, Campos, and Voorhees}]{craswell2020overview}
Nick Craswell, Bhaskar Mitra, Emine Yilmaz, Daniel Campos, and Ellen~M Voorhees. 2020.
\newblock Overview of the trec 2019 deep learning track.
\newblock \emph{arXiv preprint arXiv:2003.07820}.

\bibitem[{Datta et~al.(2024)Datta, Ganguly, MacAvaney, and Greene}]{datta2024deep}
Suchana Datta, Debasis Ganguly, Sean MacAvaney, and Derek Greene. 2024.
\newblock A deep learning approach for selective relevance feedback.
\newblock In \emph{European Conference on Information Retrieval}, pages 189--204. Springer.

\bibitem[{Datta et~al.(2022)Datta, MacAvaney, Ganguly, and Greene}]{qppdutta}
Suchana Datta, Sean MacAvaney, Debasis Ganguly, and Derek Greene. 2022.
\newblock A 'pointwise-query, listwise-document' based query performance prediction approach.
\newblock In \emph{{SIGIR}}, pages 2148--2153. {ACM}.

\bibitem[{Dhole(2023)}]{dhole2023large}
Kaustubh Dhole. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.bigpicture-1.6} {Large language models as {S}ocio{T}echnical systems}.
\newblock In \emph{Proceedings of the Big Picture Workshop}, pages 66--79, Singapore. Association for Computational Linguistics.

\bibitem[{Dhole(2024)}]{dhole-2024-kaucus}
Kaustubh Dhole. 2024.
\newblock \href {https://aclanthology.org/2024.scichat-1.5} {{KAUCUS} - knowledgeable user simulators for training large language models}.
\newblock In \emph{Proceedings of the 1st Workshop on Simulating Conversational Intelligence in Chat (SCI-CHAT 2024)}, pages 53--65, St. Julians, Malta. Association for Computational Linguistics.

\bibitem[{Dhole and Agichtein(2024{\natexlab{a}})}]{dhole2024llmjudge}
Kaustubh Dhole and Eugene Agichtein. 2024{\natexlab{a}}.
\newblock \href {https://github.com/emory-irlab/argumentation-rag/blob/main/LLM_Judges_for_Argumentation.pdf} {Llm judges for retrieval augmented argumentation}.

\bibitem[{Dhole et~al.(2024{\natexlab{a}})Dhole, Bajaj, Chandradevan, and Agichtein}]{dhole2024queryexplorer}
Kaustubh Dhole, Shivam Bajaj, Ramraj Chandradevan, and Eugene Agichtein. 2024{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2024.naacl-demo.11} {{Q}uery{E}xplorer: An interactive query generation assistant for search and exploration}.
\newblock In \emph{Proceedings of the 2024 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 3: System Demonstrations)}, pages 107--115, Mexico City, Mexico. Association for Computational Linguistics.

\bibitem[{Dhole et~al.(2023)Dhole, Gangal, Gehrmann, Gupta, Li, Mahamood, Mahadiran, Mille, Shrivastava, Tan et~al.}]{dhole2023nl}
Kaustubh Dhole, Varun Gangal, Sebastian Gehrmann, Aadesh Gupta, Zhenhao Li, Saad Mahamood, Abinaya Mahadiran, Simon Mille, Ashish Shrivastava, Samson Tan, et~al. 2023.
\newblock Nl-augmenter: A framework for task-sensitive natural language augmentation.
\newblock \emph{Northern European Journal of Language Technology}, 9(1).

\bibitem[{Dhole and Agichtein(2024{\natexlab{b}})}]{dhole2024genqrensemble}
Kaustubh~D Dhole and Eugene Agichtein. 2024{\natexlab{b}}.
\newblock Genqrensemble: Zero-shot llm ensemble prompting for generative query reformulation.
\newblock In \emph{European Conference on Information Retrieval}, pages 326--335. Springer.

\bibitem[{Dhole et~al.(2024{\natexlab{b}})Dhole, Chandradevan, and Agichtein}]{dhole2024generative}
Kaustubh~D Dhole, Ramraj Chandradevan, and Eugene Agichtein. 2024{\natexlab{b}}.
\newblock Generative query reformulation using ensemble prompting, document fusion, and relevance feedback.
\newblock \emph{arXiv preprint arXiv:2405.17658}.

\bibitem[{Khattab and Zaharia(2020)}]{khattab2020colbert}
Omar Khattab and Matei Zaharia. 2020.
\newblock Colbert: Efficient and effective passage search via contextualized late interaction over bert.
\newblock In \emph{Proceedings of the 43rd International ACM SIGIR conference on research and development in Information Retrieval}, pages 39--48.

\bibitem[{Lin et~al.(2021)Lin, Ma, Lin, Yang, Pradeep, and Nogueira}]{lin2021pyserini}
Jimmy Lin, Xueguang Ma, Sheng-Chieh Lin, Jheng-Hong Yang, Ronak Pradeep, and Rodrigo Nogueira. 2021.
\newblock Pyserini: A python toolkit for reproducible information retrieval research with sparse and dense representations.
\newblock In \emph{Proceedings of the 44th International ACM SIGIR Conference on Research and Development in Information Retrieval}, pages 2356--2362.

\bibitem[{Macdonald and Tonellotto(2020)}]{pyterrier2020ictir}
Craig Macdonald and Nicola Tonellotto. 2020.
\newblock Declarative experimentation ininformation retrieval using pyterrier.
\newblock In \emph{Proceedings of ICTIR 2020}.

\bibitem[{Meta(2024{\natexlab{a}})}]{llama3modelcard}
Meta. 2024{\natexlab{a}}.
\newblock \href {https://github.com/meta-llama/llama3/blob/main/MODEL_CARD.md} {Llama 3 model card}.

\bibitem[{Meta(2024{\natexlab{b}})}]{llama3_1modelcard}
Meta. 2024{\natexlab{b}}.
\newblock \href {https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct} {Llama 3.1 model card}.

\bibitem[{Nogueira et~al.(2020)Nogueira, Jiang, Pradeep, and Lin}]{nogueira-etal-2020-document}
Rodrigo Nogueira, Zhiying Jiang, Ronak Pradeep, and Jimmy Lin. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.findings-emnlp.63} {Document ranking with a pretrained sequence-to-sequence model}.
\newblock In \emph{Findings of the Association for Computational Linguistics: EMNLP 2020}, pages 708--718, Online. Association for Computational Linguistics.

\bibitem[{Nogueira et~al.(2019)Nogueira, Yang, Cho, and Lin}]{monobert}
Rodrigo Nogueira, Wei Yang, Kyunghyun Cho, and Jimmy Lin. 2019.
\newblock Multi-stage document ranking with bert.
\newblock \emph{arXiv preprint arXiv:1910.14424}.

\bibitem[{OpenAI(2023)}]{openai2023gpt35}
OpenAI. 2023.
\newblock \href {https://www.openai.com} {Gpt-3.5}.

\bibitem[{{OpenAI}(2024{\natexlab{a}})}]{gpt4turbo0409}
{OpenAI}. 2024{\natexlab{a}}.
\newblock \href {https://platform.openai.com/docs/models/gpt-4-turbo} {{GPT-4-turbo-0409}}.

\bibitem[{{OpenAI}(2024{\natexlab{b}})}]{gpt4omini}
{OpenAI}. 2024{\natexlab{b}}.
\newblock \href {https://platform.openai.com/docs/models/gpt-4o-mini} {{GPT-4o mini}}.

\bibitem[{Parry et~al.(2024)Parry, MacAvaney, and Ganguly}]{parry2024top}
Andrew Parry, Sean MacAvaney, and Debasis Ganguly. 2024.
\newblock Top-down partitioning for efficient list-wise ranking.
\newblock \emph{arXiv preprint arXiv:2405.14589}.

\bibitem[{Pradeep et~al.(2023{\natexlab{a}})Pradeep, Sharifymoghaddam, and Lin}]{pradeep2023rankvicuna}
Ronak Pradeep, Sahel Sharifymoghaddam, and Jimmy Lin. 2023{\natexlab{a}}.
\newblock Rankvicuna: Zero-shot listwise document reranking with open-source large language models.
\newblock \emph{arXiv preprint arXiv:2309.15088}.

\bibitem[{Pradeep et~al.(2023{\natexlab{b}})Pradeep, Sharifymoghaddam, and Lin}]{pradeep2023rankzephyr}
Ronak Pradeep, Sahel Sharifymoghaddam, and Jimmy Lin. 2023{\natexlab{b}}.
\newblock Rankzephyr: Effective and robust zero-shot listwise reranking is a breeze!
\newblock \emph{arXiv preprint arXiv:2312.02724}.

\bibitem[{Qin et~al.(2024)Qin, Jagerman, Hui, Zhuang, Wu, Yan, Shen, Liu, Liu, Metzler, Wang, and Bendersky}]{qin-etal-2024-large}
Zhen Qin, Rolf Jagerman, Kai Hui, Honglei Zhuang, Junru Wu, Le~Yan, Jiaming Shen, Tianqi Liu, Jialu Liu, Donald Metzler, Xuanhui Wang, and Michael Bendersky. 2024.
\newblock \href {https://aclanthology.org/2024.findings-naacl.97} {Large language models are effective text rankers with pairwise ranking prompting}.
\newblock In \emph{Findings of the Association for Computational Linguistics: NAACL 2024}, pages 1504--1518, Mexico City, Mexico. Association for Computational Linguistics.

\bibitem[{Robertson et~al.(1992)Robertson, Walker, Hancock-Beaulieu, Gull, and Lau}]{robertson1992okapi}
Stephen~E Robertson, Steve Walker, Micheline Hancock-Beaulieu, Aarron Gull, and Marianna Lau. 1992.
\newblock Okapi at trec.
\newblock In \emph{Text retrieval conference}, pages 21--30.

\bibitem[{Sun et~al.(2023)Sun, Yan, Ma, Wang, Ren, Chen, Yin, and Ren}]{sun2023chatgpt}
Weiwei Sun, Lingyong Yan, Xinyu Ma, Shuaiqiang Wang, Pengjie Ren, Zhumin Chen, Dawei Yin, and Zhaochun Ren. 2023.
\newblock Is chatgpt good at search? investigating large language models as re-ranking agents.
\newblock In \emph{Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing}, pages 14918--14937.

\bibitem[{Wolf et~al.(2020)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac, Rault, Louf, Funtowicz et~al.}]{wolf2020transformers}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, R{\'e}mi Louf, Morgan Funtowicz, et~al. 2020.
\newblock Transformers: State-of-the-art natural language processing.
\newblock In \emph{Proceedings of the 2020 conference on empirical methods in natural language processing: system demonstrations}, pages 38--45.

\bibitem[{Yue et~al.(2023)Yue, Rabhi, Moreira, Wang, and Oldridge}]{yue2023llamarec}
Zhenrui Yue, Sara Rabhi, Gabriel de Souza~Pereira Moreira, Dong Wang, and Even Oldridge. 2023.
\newblock Llamarec: Two-stage recommendation using large language models for ranking.
\newblock \emph{arXiv preprint arXiv:2311.02089}.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt]{article}

\usepackage{EMNLP2023}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{listings}
\usepackage{booktabs}
\usepackage{graphicx} % Needed for including graphics
\usepackage{color}
\definecolor{codegreen}{rgb}{0,0.6,0}
\definecolor{codegray}{rgb}{0.5,0.5,0.5}
\definecolor{codepurple}{rgb}{0.58,0,0.82}
\definecolor{backcolour}{rgb}{0.80, 0.92, 0.95}
\lstdefinestyle{mystyle}{
    backgroundcolor=\color{backcolour},   
    commentstyle=\color{codegreen},
    keywordstyle=\color{magenta},
    numberstyle=\tiny\color{codegray},
    stringstyle=\color{codepurple},
    basicstyle=\footnotesize,
    breakatwhitespace=false,         
    breaklines=true,                 
    captionpos=b,                    
    keepspaces=true,                 
    numbers=left,                    
    numbersep=5pt,                  
    showspaces=false,                
    showstringspaces=false,
    showtabs=false,                  
    tabsize=2
}
\title{PyTerrier-GenRank: The PyTerrier Plugin for Reranking with Large Language Models}

\author{Kaustubh D. Dhole \\
  Department of Computer Science \\
  Emory University\\
  Atlanta, GA -- USA \\
  \texttt{\textcolor{darkblue}{kdhole@emory.edu}}
 }

\begin{document}
\maketitle
\begin{abstract}
\vspace{-5pt}
Using LLMs as rerankers requires experimenting with various hyperparameters, such as prompt formats, model choice, and reformulation strategies. We introduce~\href{https://github.com/emory-irlab/pyterrier_genrank}{\underline{\textbf{PyTerrier-GenRank}}}, a PyTerrier plugin to facilitate seamless reranking experiments with LLMs, supporting popular ranking strategies like pointwise and listwise prompting. We validate our plugin through HuggingFace and OpenAI hosted endpoints.
\end{abstract}

\section{Introduction}

Generative reranking methods have recently gained popularity for their ability to produce a more optimal ranking of documents by prompting large language models. These methods are also employed in Retrieval Augmented Generation~\cite{dhole-2024-kaucus} paradigms as they are effective for enhancing generation through the incorporation of retrieved information. Typically, a sparse retriever such as BM25~\cite{robertson1992okapi} or a dense retriever like ColBERT~\cite{khattab2020colbert} is utilized to retrieve a small subset (ranging from 20 to 1000) from among millions or billions of text pieces. This subset is then forwarded to a second-stage generative reranker for reranking.

The increasing adoption of such reranking methods necessitates rigorous testing of various hyperparameters, including employing different prompts, evaluating performance with different model weights, and using it in conjunction with various retrieval, reranking, and reformulation~\cite{dhole2024queryexplorer, dhole2024generative} pipelines.

In that regard, we introduce --~\textbf{PyTerrier-GenRank} -- to encourage rapid experimentation of retrieval pipelines employing recently popular LLM-based reranking methods. The code includes PyTerrier~\cite{pyterrier2020ictir} wrappers over many helper functions of the RankLLM repository~\cite{pradeep2023rankvicuna, pradeep2023rankzephyr}. Our plugin can be quickly installed in any Python environment and is available~\href{https://github.com/emory-irlab/pyterrier_genrank}{\textcolor{codegreen}{at this link}}\footnote{\href{https://github.com/emory-irlab/pyterrier_genrank}{\textcolor{codegreen}{github.com/emory-irlab/pyterrier\_genrank}}}.

\label{Listing1}
\begin{lstlisting}[language=Python, caption=Retrieving using a listwise reranker]
 
 import pyterrier as pt
 from rerank import LLMReRanker

 bm25 = pt.BatchRetrieve(dataset.get_index(), wmodel="BM25")

 reranker = LLMReRanker("castorini/rank_vicuna_7b_v1")

 pipeline = bm25 % 100 
    >> pt.text.get_text(index, 'text') 
    >> reranker

 pipeline.search('Indian restaurants')
 
\end{lstlisting}
\section{Generative Reranking}
Large Language Models (LLMs) have been widely used for reranking tasks in various ways involving novel hyperparameters. Among the most popular ones have been ranking documents in a pointwise, pairwise and listwise fashion. MonoBERT~\cite{monobert} and MonoT5~\cite{nogueira-etal-2020-document} are among the earliest pointwise ranking methods that were trained to emit a single scalar relevance score for a query-document pair. In contrast, pairwise approaches~\cite{qin-etal-2024-large} involved prompting models with a query and two documents to compare and rank them.~\texttt{RankVicuna}~\cite{pradeep2023rankvicuna},~\texttt{RankZephyr}~\cite{pradeep2023rankzephyr}, and~\texttt{RankGPT}~\cite{sun2023chatgpt, openai2023gpt35} have explored listwise ranking strategies by prompting their respective models with a list of documents and generating a ranked list of document IDs. These paradigms have also been investigated in the context of recommender systems~\cite{yue2023llamarec}.

The number of hyperparameters involved in creating an optimal retriever, the choice of prompts, the number of documents, the weight of the reformulation strategy, etc., all influence the final ranking set in various ways. Besides, the increasing number of LLMs released by various organizations further necessitates rigorous testing to evaluate performance across different settings. To ensure the validity and reproducibility of results, it is imperative that researchers and developers can easily replicate experiments across these novel sets of hyperparameters. 

\section{Python IR Toolkits}
We choose PyTerrier~\cite{pyterrier2020ictir} due to its modularity and Python-friendly interface. It is built on the Terrier search engine and excels in modular, rapid experimentation for IR research, allowing users to easily integrate and evaluate various retrieval and re-ranking models. It has been employed in various studies to evaluate retrieval effectiveness~\cite{dhole2024genqrensemble, datta2024deep, parry2024top, qppdutta, dhole2024llmjudge}.
\begin{table}[h!]
\centering
\resizebox{\columnwidth}{!}{
%\begin{tabular}{>{\raggedright\arraybackslash}p{1.2cm} >{\raggedright\arraybackslash}p{2cm} >{\raggedright\arraybackslash}p{2cm}}
\begin{tabular}{llc}
\toprule
\hline
\textbf{Training} & \textbf{Name} & \textbf{nDCG@10} \\
\hline
\midrule
0-shot & BM25 & .480 \\
0-shot & gpt-35-turbo$^{\beta}$~\cite{openai2023gpt35} & .660 \\
0-shot & gpt-4-turbo-0409$^{\beta}$~\cite{gpt4turbo0409} & .701 \\
0-shot & gpt-4o-mini~\cite{gpt4omini} & \textbf{.710} \\
0-shot & Llama-3-8B-Instruct~\cite{llama3modelcard} & .572 \\
0-shot & Llama-3.1-8B-Instruct~\cite{llama3_1modelcard} & .594 \\
0-shot & Llama-Spark (8B)~\cite{spark} & \textbf{.612} \\
\hline
Trained & Rank-Vicuna$^{\beta}$~\cite{pradeep2023rankvicuna} & .672 \\
Trained & Rank-Zephyr$^{\beta}$~\cite{pradeep2023rankzephyr} & .711 \\
\hline
\bottomrule
\end{tabular}
}
\caption{Comparing reranking performance of different models for TREC-DL 2019~\cite{craswell2020overview} using PyTerrier\_Genrank}
\label{tableres}
\end{table}

\section{Generative Reranking Wrappers}
Our plugin employs PyTerrier wrappers over reranking class functions modified from the Pyserini~\cite{lin2021pyserini} based~\href{https://github.com/sunnweiwei/RankGPT}{RankGPT} and~\href{https://github.com/castorini/rank_llm}{RankLLM} repositories. We allow additional hyperparameters to be expressed from function calls. We provide sample code to execute HuggingFace~\cite{wolf2020transformers} models and OpenAI-based endpoints.

\section{Sanity Check}
As a sanity check, we re-evaluate some previously published rerankers like RankVicuna, RankGPT, using this framework. Those marked with superscript $^{\beta}$ have already been previously tested with the Pyserini framework in their respective papers. We also evaluate the zero-shot retrieval effectiveness of newly published instruction-tuned models like LLama-Spark.

We find that among the 8B zero-shot approaches, LLama-Spark is the most effective as a reranker. The complete results are shown in Table~\ref{tableres}.

\section*{Ethics Statement}
Large Language Models are embedded within a broader socio-technical framework~\cite{dhole2023large,dhole2023nl}. When deploying them for tasks such as ranking and recommendations, it's crucial to assess their long-term impacts, including the potential for creating echo chambers and the risk of generating biased suggestions.

% Entries for the entire Anthology, followed by custom entries
\bibliography{anthology,custom}
\bibliographystyle{acl_natbib}

\appendix

\end{document}